# MonoSOWA: Scalable monocular 3D Object detector Without human Annotations

### v1

@@ IV-B Ablations
TABLE IV: Canonical Object Space (COS) ablation on KITTI validation set under full supervision from human labels.
Canonical Object Space
As seen in table IV, disabling COS and training on both datasets jointly decreases the average precision of both BEV and 3D. It may be surprising that when training on both KITTI+K360, the model is not broken completely – we speculate that the network is able to learn to classify from which of the two datasets the image comes from and adjusts its outputs accordingly; such an approach obviously does not generalize well. Note that only in this ablation, we use human labels to eliminate any effects of pseudo-labels.
Number of frames used in aggregation
No. of APBEV/AP3D@0.5
frames Easy Moderate Hard
$\pm$ 1 53.26/49.35 41.34/35.86 34.77/29.14
$\pm$ 10 59.76/61.55 44.08/37.09 36.99/33.15
$\pm$ 30 59.22/51.18 42.65/36.64 35.91/29.77
TABLE V: Number of frames used in the aggregation process ablation on KITTI validation set.
table V shows the ablation study on the number of frames used in the aggregation process. Perhaps surprisingly, aggregating over longer sequences does not yield better performance. The aggregation can recover from the error caused by Metric3D [3, 9] as in fig. 4. We speculate that the distant car poses a significant challenge for the detector itself, therefore creating precise 3D labels for such objects does not translate to better 3D detection accuracy.
APBEV/AP3D@0.5
$\mathbf{\alpha}$ Easy Moderate Hard
1 53.02/48.09 41.29/34.92 34.48/28.66
5 55.14/50.83 42.46/36.59 35.92/29.85
10 56.56/48.90 41.32/35.00 34.53/28.76
15 55.78/51.86 42.48/36.63 35.72/29.81
TABLE VI: Steepness parameter $\alpha$ of the Saturated Closeness Criterion ablation on KITTI validation set.
Steepness parameter in Saturated Closeness Criterion
In this ablation, we explore possible values for the steepness parameter $\alpha$ in the Saturated Closeness Criterion, which controls what pseudo-LIDAR points are discarded as outliers. As shown in table VI, $\alpha$ equal to 10 achieves the best performance.
@@ V Conclusion
A novel method for training monocular 3D detectors without domain-specific human annotations was presented. The method exploits temporal consistency in video sequences to automatically create 3D labels of objects (cars) and, as such, does not require human labels or additional sensors such as LiDAR. The method is able to compensate for camera focal length differences to significantly improve its scalability across multiple datasets, and furthermore, it works out of the box in previously unseen camera setups.
The method is evaluated on two large-scale public datasets, where despite using no human labels, it outperforms prior work by a significant margin both for BEV and 3D when using the stricter 0.5 IoU evaluation. Last but not least, the method shows its versatility by also being a powerful pre-training method for fully-supervised monocular 3D detection, where the final accuracy increases by approximately 6 percent points over traditional object detector.
The main limitation is detecting objects far away, which is not primarily due to the auto-labelling procedure but due to the inherent limitation of the trained monocular detector to infer the distance of objects which are only a couple of pixels high.

### v2

@@ IV-B Ablations
In this section, we present two ablation experiments we deemed the most important, and, due to lack of space, we kindly refer the reader to the Supplementary material for further ablations. The auto-labelling pipeline ablations use KITTI [10] training set for evaluation as it best measures the quality of the auto-generated labels, while the ablations involving a trained model (MonoDETR) use the respective KITTI training/validation sets for training and evaluation.
First, we inspect the impact of the proposed Local Object Motion Model and Saturated Closeness Criterion. As seen in  table VII, both components are crucial to achieve the best accuracy of generated pseudo-labels.
LOMM SCC APBEV/AP3D@0.5
Easy Moderate Hard
20.41/17.40 18.22/15.88 19.99/13.75
20.37/17.16 18.26/15.76 20.04/13.66
35.89/29.05 26.32/17.43 28.87/19.19
39.22/32.29 33.00/23.14 31.09/21.34
TABLE VII: Local Object Motion Model (LOMM) and Saturated Closeness Criterion (SCC) ablation on the KITTI [10] training set.
In the second ablation, we don’t use Canonical Object Space and instead naively combine KITTI and KITTI-360 datasets. As expected, the average precision decreases when KITTI-360 data are appended to KITTI (see table VIII), because the network focuses on camera setup of the larger K360 dataset, and it is only with COS enabled that the network actually benefits from the added data. Note that only in this ablation, we use human labels to eliminate any effects of pseudo-labels.
@@
TABLE VIII: Canonical Object Space (COS) ablation using human labels, evaluated on KITTI validation set.
@@ V Conclusion
A novel method for training monocular 3D detectors without domain-specific human annotations was presented. The method exploits temporal consistency in video sequences to automatically create 3D labels of objects (cars) and, as such, does not require human labels or additional sensors such as LiDAR. The method uses newly proposed Local Object Motion Model (LOMM) to disentangle object movement source between subsequent frames, is approximately 700 times faster than current state of the art and is able to compensate for camera focal length differences to improve scalability across multiple datasets.
The method is evaluated on three large-scale public datasets, where despite using no human labels, it outperforms prior work by a significant margin both for BEV and 3D when using the stricter 0.5 IoU evaluation. The method also shows its versatility by being a powerful pre-training tool for fully-supervised monocular 3D detection, and last but not least, shows that combining pseudo-labels from multiple datasets can achieve comparable accuracy to using human labels from a single dataset. The main limitation is detecting objects far away, which is not primarily due to the auto-labelling procedure but due to the inherent limitation of the trained monocular detector to infer the distance of objects which are only a couple of pixels high.